Westerville Company reported the following results from last year’s operations: Sales $ 1,200,000 Variable expenses 320,000 Contribution margin 880,000 Fixed expenses 640,000 Net operating income $ 240,000 Average operating assets $ 600,000 At the beginning of this year, the company has a $150,000 investment opportunity with the following cost and revenue characteristics: Sales $ 240,000 Contribution margin ratio 50 % of sales Fixed expenses $ 84,000 The company’s minimum required rate of return is 15%.

10-a. If Westerville’s chief executive officer will earn a bonus only if her ROI from this year exceeds her ROI from last year, would she pursue the investment opportunity?

  • Yes

  • No

10-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No

11.  What is last year’s residual income?

12. What is the residual income of this year’s investment opportunity?

13. If the company pursues the investment opportunity and otherwise performs the same as last year, what residual income will it earn this year?
Answer #1
Part-1
Net Operating Income from new Investment Opportunity = Contribution Margin- Fixed Expense
( 240000X 50%)-84000=$36000
Existing Net Operating Income    = $240000
New Net Operating Inocome = 360000+240000=276000
Existing ROI= 240000/600000=40%
New ROI= 276000/750000=36.8%
No, he will not persued this opportunity
Part-2
Residual Income = Net Operating Income - ( Average Investment X Required Return)
240000-(600000X 15%)=$150000
Part-3
Residual Income = Net Operating Income - ( Average Investment X Required Return)
276000-(750000*15%)=$163500